OpenStudy (anonymous):

Choose the equation of the line with m=9, y-intercept (0,-2)

OpenStudy (nenadmatematika):

y+2=9(x-0) so y=9x-2

OpenStudy (anonymous):

y=mx+b You're given m and b, so just substitute them here: y=9x+(-2)=9x-2